OpenStudy (linn99123):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 3m - 5 = 7m - 21 @biby @Squirrels

OpenStudy (linn99123):

@bibby

OpenStudy (cj49):

3m-5=7m-21 3m-7m=-21+5 -4m=-16 m=4

OpenStudy (linn99123):

you could have explaind..... but its fine

OpenStudy (bibby):

the explanation is the same among all equations. isolate what you're solving for.

OpenStudy (linn99123):

bibby if i dye my hair brown will i be smart like you?

OpenStudy (bibby):

I'm balding. Does that mean I'm becoming stupider as the days go by?

OpenStudy (linn99123):

yes sir ;-;

OpenStudy (cj49):

3m-5=7m-21(subtract 7m from both the sides) -4m-5=-21(add 5 to both the sides) -4m=-16(divide both the sides by -4) m=16

OpenStudy (linn99123):

Thank yew c: And bibby ill get a foot photo soon ... ive been busy hun